### v2

# Learning Lens Blur Fields

Abstract
Optical blur is an inherent property of any lens system and is challenging to model in modern cameras because of their complex optical elements. To tackle this challenge, we introduce a high-dimensional neural representation of blur—the lens blur field—and a practical method for acquiring it. The lens blur field is a multilayer perceptron (MLP) designed to (1) accurately capture variations of the lens 2D point spread function over image plane location, focus setting and, optionally, depth and (2) represent these variations parametrically as a single, sensor-specific function. The representation models the combined effects of defocus, diffraction, aberration, and accounts for sensor features such as pixel color filters and pixel-specific micro-lenses. To learn the real-world blur field of a given device, we formulate a generalized non-blind deconvolution problem that directly optimizes the MLP weights using a small set of focal stacks as the only input. We also provide a first-of-its-kind dataset of 5D blur fields—for smartphone cameras, camera bodies equipped with a variety of lenses, etc. Lastly, we show that acquired 5D blur fields are expressive and accurate enough to reveal, for the first time, differences in optical behavior of smartphone devices of the same make and model. Code and data can be found at blur-fields.github.io.
Index Terms:
Modeling Point Spread Functions, Computational Photography, Neural Representations
Figure 1: Illustration of learning lens blur fields. We prescribe an easy-to-use method for acquiring a lens blur field—a neural representation that captures high-dimensional variations in the point spread function of an imaging system. (A) We first capture focal stacks of calibration patterns displayed on a monitor, for example, using the Google Pixel 4 as depicted. This particular smartphone camera has a dual-pixel sensor which places two pixels behind a microlens focused on the objective lens. (B-C) We use captured calibration patterns to optimize a multi-layer perceptron that parameterizes variations in the PSF ($u$, $v$) across image plane positions ($x$, $y$) and focus ($f$). We visualize slices of the recovered 5D blur field for the left-looking and right-looking green-channel pixels of the sensor.
I Introduction
Accurately modeling an optical imaging system requires characterizing the blur induced by optical aberration, defocus, and diffraction. Such models of optical blur underlie many computational imaging [1, 2, 3] and sensing  [4, 5, 6] techniques and are becoming increasingly important for high-performance computer vision in automotive [7] and mobile photography applications [8, 9, 10, 11, 12].
Despite the potentially wide-ranging impact of such models, existing approaches for estimating optical blur in real-world camera systems trade-off representational accuracy against ease of acquisition. Accurately representing optical blur across multiple dimensions requires densely sampling the parameter space, which increases acquisition time and estimation complexity. To address this major limitation, we propose a novel method for blur acquisition that enables highly accurate modeling of smartphone and interchangeable-lens camera systems without sacrificing scalability across many potential dimensions (2D pixel location, depth, focus setting, wavelength, etc.) or ease of use. A key property of our approach is that it outputs a model that is continuous, making it possible to represent both large scale and minute blur variations across all those dimensions. Although our specific focus in this paper is conventional cameras for which a thin-lens model can serve as a (very crude) approximation, generalization to systems employing coded optics [13, 14, 15, 16, 17] may be possible as well.
I-A Background and Related Work
Optical blur is typically characterized using a point spread function (PSF), which describes the image formed by a point light source on the camera’s sensor plane. Recovering the PSF of a camera is a challenging problem: blur can vary considerably over the sensor plane and is dependent on the distance to the light source, the camera’s focus setting, aperture, wavelength, etc.
Historically, PSFs have been modeled in three different ways: parametrically, non-parametrically, or via geometric or Fourier optics simulation. Parametric methods [18, 10, 19, 20, 21] rely on idealized models of lenses and image formation and thus have limited expressiveness. For example, models based on Gaussian kernels [22, 23] or hand-crafted formulations [24, 25] are too simple to model real-world camera PSFs; Zernike polynomials assume rotational symmetry and circular pupils [26, 18, 20, 27, 28, 29, 30]; and Seidel aberrations [20, 19, 21] assume axially-symmetric lens systems and cannot model blur of dual-pixel auto-focus systems, which do not have such symmetry.
Non-parametric methods sample the space of a camera’s PSFs by acquiring a sparse grid of 2D PSFs over the sensor plane [31, 32, 22, 33] and optionally over focus settings [31] and depth [32]. These approaches are fundamentally limited by a density-accuracy tradeoff because they solve an independent, local blind [31, 24, 34, 35, 36, 37] or non-blind [38, 39, 40, 41, 42] deconvolution problem for each PSF, which becomes degenerate as the image patches used for deconvolution become smaller and/or the PSFs spatial extent becomes larger [32, 43]. As a result, these methods sample PSFs sparsely, limiting their ability to model PSFs that vary rapidly along one or more dimensions [44]. Although in theory these limitations can be side-stepped by adhering to the PSF definition and imaging a point light source to capture the PSF directly, in practice it is difficult to construct an ideal point light source [45], i.e., a source with a perfectly circular shape that projects to a region smaller than a pixel and has uniform radiant intensity. Moreover, sources with the micron-scale sizes required for the tiny pixels in modern smartphones are subject to diffraction [46, 47] and have very low light efficiency, making it even more difficult to treat their images as that of a “ground truth PSF”. For all these reasons, accurate estimation of spatially varying PSFs has long relied on deconvolution of known patterns [31] as opposed to direct imaging.
Optical simulation systems [48, 49, 50, 51, 52], on the other hand, assume that a physical model of the complete camera-plus-lens system is known precisely and that the corresponding real-world devices conform to it exactly. In many cases, however, precise models of widely-used compound lens systems may be proprietary or unknown [53], and their precision cannot be taken for granted due to device age, mechanical impacts from daily use, manufacturing variabilities, etc. While many deep learning techniques tackle the blind [54, 55, 56, 57, 58, 59] or non-blind [60, 61, 62, 63, 64] deconvolution problems, none of these methods apply to the high-dimensional, spatially-varying cases we consider. Moreover, these methods rely on a discrete PSF representation instead of an MLP and thus suffer from the curse of dimensionality as PSF dimensions increase. Consequently, the recovery of accurate, densely-sampled, high-dimensional PSFs for commodity cameras has remained beyond the reach of the broader computational photography community due to these technical obstacles.
I-B Contributions
In this paper, we seek to fill this gap by introducing a novel neural representation of blur—the lens blur field—along with a practical method of acquiring it. The lens blur field represents a camera’s PSF as a multi-layer perceptron (MLP) that is designed to capture blur variations across sensor position, focus setting and, optionally, depth. This representation is particularly suitable for PSF acquisition because it yields a continuous representation of the PSF across multiple dimensions instead of sparse discrete samples; it can be optimized by leveraging the advanced MLP training schemes already available; it can learn high-dimensional signals more accurately while using only a fraction of the parameters of non-parametric approaches [65, 66]; and it can accurately capture difficult-to-model effects such as defocus, chromatic aberration, and higher-order aberrations with relatively few parameters. Overall, the lens blur field diminishes the tradeoff between representational accuracy and ease of acquisition by leveraging the interpolation properties of MLPs [67, 68], enabling scalable modeling within the input dimensions while accurately capturing detailed optical lens effects.
Recovering the lens blur field involves training its MLP on a dataset of focal stacks of a computer screen that can be easily acquired in less than five minutes. Our approach is applicable to a wide array of devices, from smartphone cameras and dual-pixel sensors to single-lens reflex (SLR) cameras and lenses and we show how to reconstruct 5D and 6D blur fields from such data.
Our method offers the following key technical contributions over the state of the art:
  • We develop a PSF acquisition procedure that enables reliable, fast, lightweight acquisition of continuous, high-dimensional PSFs of commodity cameras.
  • We reduce the problem of high-dimensional deconvolution to the problem of training a continuous MLP, whose network architecture can remain fixed across all devices we consider (smartphones, SLR lenses).
  • We assemble a first-of-its-kind dataset of 5D and 6D lens blur fields.
Our work provides new insights and enables new applications. For example we show that PSFs vary considerably between smartphone models, including the iPhone 12 Pro and iPhone 14 Pro; that dual-pixel auto-focus systems have PSFs that depart significantly from the analytical models used hitherto; and that lens blur fields are reliable and expressive enough to reveal subtle differences in the optical behavior of smartphone devices of the same make and model, yielding a device-specific “blur signature”. We also show that lens blur fields can be used as space-efficient computational proxies of device optics, allowing us to apply high-dimensional blurs to render images with device-specific PSFs (Fig. 8). Taken together, we provide new ways to both model and simulate the performance of common imaging systems, including smartphone cameras used by billions worldwide.
Figure 2: Overview of learning lens blur fields. (A) Our approach learns a blur field by solving a non-blind deconvolution problem between blur-free calibration patterns $\hat{I}$ and blurry images $I$ of the patterns. Blur-free images are computed by applying perspective projection, lens distortion, and radiometric compensation to the albedo $A$ at the scene plane (top left). (B) An MLP is used to represent the lens blur field, which parameterizes the PSF over sensor position $\mathbf{x}$, target distance $d_{i}$, focus distance $f_{i}$, and a displacement $\mathbf{u}$ from $\mathbf{x}$. We train the MLP on a dataset of blur-free and captured focal stack images of synthetic noise patterns [69] and optimize MLP parameters to minimize the error between predicted and captured blurry image patches. Choice of noise patterns are discussed in Supplement Sec. S-IV.
II Learning Lens Blur Fields
Our approach leverages multilayer perceptrons (MLPs) both for modeling the high-dimensional PSF of a camera system and for acquiring it from image data.
II-A Lens Blur Field Representation
Consider an incoherent point light source $\mathbf{p}\in\mathbb{R}^{3}$ at distance $d$ from the front pupil plane of a lens (see Fig. 2). Let $\mathbf{\tilde{x}}\in\mathbb{R}^{2}$ be the perspective projection of point $\mathbf{p}$ onto the camera’s sensor plane and let $\mathbf{x}\in\mathbb{R}^{2}$ be the point’s actual projection after accounting for non-linear geometric distortions caused by the lens. Since a camera’s focus setting $f$ may affect the overall magnification of captured images [70] and since dual-pixel AF sensors induce disparity between left-looking and right-looking pixels, in general $\mathbf{\tilde{x}}$ and $\mathbf{x}$ depend on both $f$ and the pixel type $c$. In the following, we express the one-to-one relation between geometrically-distorted points on the sensor plane and point sources at distance $d$ using
$\mathbf{p}~{}=~{}\mathrm{invproj}(\mathbf{x},d,f,c).$ (1)
The lens blur field, denoted by $\text{PSF}_{\theta}(\mathbf{x},d,f,\mathbf{u})$, is an MLP with trainable parameters $\theta$ that describes the blurry 2D image of a light source at $\mathbf{p}$ as a function of the displacement $\mathbf{u}\in\mathbb{R}^{2}$ from $\mathbf{x}$.
For camera sensors with $C>1$ types of pixels we consider the MLP to be a vector-valued function with $C$ output dimensions. For example, $C=4$ for a sensor with a conventional RGB Bayer color filter array, $C=2$ for a monochrome sensor with dual-pixel autofocus [12] and $C=8$ for a color dual-pixel AF sensor.
II-B Continuous Image Formation
Now suppose we image a fronto-parallel plane at distance $d$ from the lens. We express the plane’s blur-free image for pixel type $c$ and focus setting $f$ as the product
$\hat{I}^{(c)}(\mathbf{x},d,f)~{}=~{}A^{(c)}(\mathbf{p})~{}E^{(c)}(\mathbf{p},f){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},}$ (2)
where $A^{(c)}(\mathbf{p})\in[0,1]$ denotes the plane’s unit-less albedo at point $\mathbf{p}$ and $E^{(c)}(\mathbf{p},f)$ is the total irradiance incident on the sensor plane due to light falling onto (or emitted from) point $\mathbf{p}$. Intuitively, the albedo $A^{(c)}(\mathbf{p})$ represents the plane’s spatially-varying texture (or emission pattern) whereas the factor $E^{(c)}(\mathbf{p},f)$ accounts for all camera- and plane-specific radiometric quantities such as vignetting, cosine-fourth fall-off [71], the plane’s reflectance or emission properties, and the solid angle at $\mathbf{p}$ subtended by the lens pupil [72].
The plane’s blurry image is then given by the integral
$I^{(c)}(\mathbf{x},d,f)=\int_{\mathcal{U}}\text{PSF}^{(c)}_{\theta}(\mathbf{x},d,f,\mathbf{u})\cdot\hat{I}^{(c)}(\mathbf{x}-\mathbf{u},d,f)\,\mathrm{d}\mathbf{u},$ (3)
where $\mathbf{u}$ is a displacement from the projection of each point $\mathbf{x}$ on the sensor plane. Integration is performed over $\mathcal{U}$, which is the spatial support of the blur field over $\mathbf{u}$.
Learning the Blur Field
As illustrated in Fig. 2, we optimize the parameters of the blur field MLP by minimizing a photometric loss between a set of captured blurry images of known 2D patterns and corresponding rendered images estimated using Eq. 3.
Specifically, we solve the following optimization problem.
$&\operatorname*{argmin}_{\theta}\sum_{\begin{subarray}{c}t,c,i\\ \mathbf{x}\in\mathcal{P}\end{subarray}}\Bigl{\lVert}I^{(c)}_{t}(\mathbf{x},d_{i},f_{i})-\sum_{\mathbf{u}}\text{PSF}^{(c)}_{\theta}(\mathbf{x},d_{i},f_{i},\mathbf{u})\cdot\hat{I}_{t}^{(c)}(\mathbf{x}-\mathbf{u},d_{i},f_{i})\Bigr{\rVert}^{2}_{2},\\ &\text{subject to }\text{PSF}_{\theta}\geq\mathbf{0}\ \ .\end{aligned}$
The outer summation is performed over all pixel channels $c$, all sensor points $\mathbf{x}$, focus and distance settings $f_{i}$ and $d_{i}$, $1\leq i\leq N_{I}$, and for each of $N_{T}$ known patterns ($1\leq t\leq N_{T}$). The inner summation is an approximation of the integral formulation of Eq. 3 with a discrete sum over $\mathbf{u}$. We consider two instances of the training procedure depending on whether or not the distance $d_{i}$ is the same for all input images. In the former case $d$ is fixed, yielding a 5D blur field; otherwise we optimize for a 6D blur field.
Implementation Details
We train on raw captured images, with the summation over $\mathbf{x}$ in Sec. II-B taken only over pixels for which we have image measurements for a given channel. For example, for a sensor with a Bayer color filter array, we supervise the red, green, and blue channels only on coordinates $\mathbf{x}$ that are explicitly measured.
The MLP uses seven fully-connected layers of 512 channels each with ReLU activation functions. The final layer has a sigmoid activation to satisfy the non-negativity constraint. Unlike recent work employing MLPs, we do not apply positional encoding to the MLP’s input coordinates [73] as we find it unnecessary for accurate PSF acquisition. Blur fields are trained for between $4\times 10^{6}$ and $10\times 10^{6}$ iterations with the Adam optimizer ($\beta_{1}=0.5$, $\beta_{2}=0.999$) on a single NVIDIA A6000 using the tiny-cuda-nn framework [74]. Details on other hyperparameters (learning rate, exponential decay, etc.) are included in Supplement Sec. S-III. Ablations on MLP architecture are included in Supplement Sec. S-IV.
Additionally, to improve computational efficiency, we assume that the blur field is locally stationary and neglect variations over sufficiently small image patches $\mathcal{P}$ centered at $\mathbf{x}$. This allows us to (1) reduce the number of computationally-expensive MLP forward passes to a single forward pass per patch and (2) formulate the inner summation as a matrix-vector product that can be efficiently parallelized on GPUs. Note that despite employing a patch-based forward model, our approach does not suffer from the density-accuracy tradeoff characteristic of local deconvolution methods [75]. These methods rely on patches to solve a local inverse problem, thereby restricting the pixels contributing to its solution and reducing the solution’s signal-to-noise ratio. Here, deconvolution is performed globally and implicitly by the MLP, with all pixels across all images potentially contributing to it.
During training, we set one batch element to comprise a 2D PSF obtained by first randomly sampling a focus setting $f_{i}$ and a point $\mathbf{x}$ at a pixel center. Then we sample a grid of points in the two dimensions of $\mathbf{u}$ within the patch around $\mathbf{x}$ and query the MLP at these points to retrieve a 2D PSF. Sec. II-B is computed using the output of the convolution of the 2D PSF with $\hat{I}$ within the area of the patch. In our experiments, we use a batch size of one and allow 2D PSFs up to $120\times 120$ samples in size. The size of a patch $\mathcal{P}$ is set to be 1.5 times the size of the 2D PSF along each dimension, and the convolution is performed within the valid region of the patch to avoid boundary artifacts.
III Image Acquisition Procedure
We describe the procedure to acquire the blur-free and captured blurry images of the calibration patterns used to optimize the lens blur field. The steps of the procedure are outlined in Fig. 2 (A) and consist of first computing the blur-free calibration image by applying (1) perspective projection via homography, (2) lens distortion, and (3) radiometric compensation to the scene-plane albedo $A$.
III-A Homography Estimation
The perspective projection of the scene-plane albedo $A(\mathbf{p})$ onto the sensor plane (i.e., $\mathbf{p}\rightarrow\mathbf{\tilde{x}}$—see Fig. 2) is computed via a homography $H(\mathbf{\tilde{x}},d_{i},c)$. To estimate the homography, we require a pattern that enables robust point localization with sub-pixel accuracy, even in the presence of large defocus. To this end, we use two dot-grid calibration patterns consisting of white circles on a black background and black circles on a white background (shown, right). We compute the albedo dot-grid images and capture in-focus dot-grid images for which we also remove radial distortion using the procedure described in the next subsection. Following Scharstein and Szeliski [76], we binarize the dot-grid images and detect the pixel coordinates of the dot centers in each image. Then we refine the dot center locations in both images using a center-of-mass calculation on the dots in the original albedo and captured images. Finally, we compute the homography using RANSAC [77] between the dot center point correspondences in the images. Applying the homography to the scene-plane albedo as $A(H^{-1}(\mathbf{\tilde{x}},d_{i},c))$ gives the sensor-plane albedo.
III-B Lens Distortion Estimation
Now, we wish to map the sensor-plane albedo to an image of albedo that includes the effects of magnification due to changes in focus setting and the effects of lens distortion. To account for magnification changes, we acquire an undistorted captured focal stack of dot-grid images for each distance $d_{i}$ and solve for the scale transformations $S(\mathbf{\tilde{x}},d_{i},f_{i},c)$ that align the dot centers from each out-of-focus image to the single in-focus image. We apply the resulting scale transformations (one for each focus setting and distance) to the sensor-plane albedo image computed after homography estimation.
To estimate the lens distortion we use the sixth-order Brown-Conrady model of radial distortion implemented in OpenCV’s calibrateCamera function [78] and apply it to the dot-grid center positions of the captured in-focus image described in the homography estimation step. We apply this model of radial distortion $D(\mathbf{\tilde{x}},c)$ to all scaled sensor-plane albedo images over $f_{i}$ and $d_{i}$. The composition of homography $H$, scale transformation $S$, and radial distortion $D$ define the invproj operator introduced in Eq. 1, which maps points $\mathbf{x}$ on the sensor plane to points $\mathbf{p}$ on the scene plane.
III-C Radiometric Compensation
Next, we apply radiometric compensation to the distorted scene-plane albedo computed for a given $d_{i}$ and $f_{i}$. To do this, we capture two images, $I_{\text{min}}$ and $I_{\text{max}}$ under an all-black and an all-white pattern, corresponding to albedos $A(\mathbf{p})=A_{\text{min}}>0$ and $A(\mathbf{p})=A_{\text{max}}$, respectively. These images can be thought of as scaled versions of the total irradiance $E(\mathbf{x},d_{i},f_{i})$ because this factor is nearly constant over the monitor’s plane and thus it is not affected by the lens blur even when the 2D PSF at $\mathbf{x}$ has a very large spatial extent:
$I_{\text{min}}(\mathbf{x})&=A_{\text{min}}\int_{\mathbf{u}}\text{PSF}_{\theta}(\mathbf{x},d_{i},f_{i},\mathbf{u})\cdot E(\text{invproj}(\mathbf{x},d_{i},f_{i})-\mathbf{u})\mathrm{d}\mathbf{u}\\ &\approx A_{\text{min}}\cdot E(\text{invproj}(\mathbf{x},d_{i},f_{i})).\end{aligned}$
By combining Eq. 4 and its analogue for $I_{\text{max}}$ which follows trivially, the blur-free image $\hat{I}$ can now be expressed as an affine combination of $I_{\text{min}}$ and $I_{\text{max}}$.
$\hat{I}(\mathbf{x})$ $=\left[1-A(\text{invproj}(\mathbf{x},d_{i},f_{i}))\right]\cdot I_{\text{min}}(\mathbf{x})$ (5)
$+A(\text{invproj}(\mathbf{x},d_{i},f_{i}))\cdot I_{\text{max}}(\mathbf{x}).$
III-D Image Capture
Finally, we describe the procedure for capturing focal stack images displayed on a monitor. We use a 32-inch 4K LED monitor to display the calibration patterns for both smartphone and SLR captures. We mount each camera on a tripod and position it fronto-parallel to the monitor. Depending on the distance $d$ to the scene plane, the image of the calibration pattern may span the entire sensor plane or just a portion of it. In the former case, a single focal stack of images is captured for lens distortion estimation and radiometric compensation. In the latter, the camera and/or monitor are physically moved to maintain $d$ while also adjusting the patterns to project to different parts of the sensor plane (see result in Sec. IV). For smartphone experiments, focal stacks are automatically captured using software, and we sample a range of focus positions denoted as $N_{I}\in\{7,24\}$, which depends on hardware and device limitations. Finally, we set the ISO to the lowest possible value and capture bursts with exposure times of 100 ms (for iPhones) or 250 ms (for Pixel and SLR) to avoid capturing screen refresh, PWM, and flicker. Supplement Sec. S-II contains additional details on the acquisition procedure for all devices in our dataset.
Figure 3: Simulated recovery of spatially invariant dual-pixel blur kernels created with the parametric models proposed by Punnappurath et al. [10]. Each sub-panel gives a side-by-side comparison of the ground truth (left), our recovered PSF (middle left), the result of Joshi et al. [31] (middle right), and Mannan & Langer [32] (right). Root-mean-square error (RMSE) is estimated on the inset regions shown in Fig. 3 to mitigate bias from large regions of zero-valued pixels. The values are shown in the bottom left corner of each plot. Supplement Sec. S-IV contains more extensive comparisons and cross-sectional plots.
IV Results
We evaluate our method in simulation and on captured data, and perform quantitative and qualitative comparisons to other non-blind PSF estimation methods. We also describe recovered blur fields for a wide range of lenses and devices, ranging from iPhone and Google Pixel cameras to varying SLR lenses.
IV-A Simulated Evaluation
To evaluate our method in simulation, we use calibration patterns blurred with a known PSF. Specifically, we use the synthetic noise patterns [69] shown in Fig. 2 and blur them with a spatially-invariant parametric model of lens blur for dual-pixel cameras [10]. The focus is adjusted by varying the radius of the PSFs using the thin lens model [79], with a 30 mm focal length, 40 cm calibration target distance, and linear sampling in diopter space. The maximum and minimum defocus blur radii are set to 24 and 0.5 pixels respectively. The resulting focal stack comprises $N_{I}=21$ images for each synthetic noise pattern. Gaussian read noise with a standard deviation of 0.01 is added to the blurred images. Additionally, we simulate a 16-bit capture with Poisson noise. Overall, our noise model incorporates read noise (Gaussian), shot noise (Poisson), and ADC (quantization) noise.
We train on the paired blur-free and blurry images, holding out every other image in the focal stack for validation. The proposed approach is compared to PSFs reconstructed using the methods of Joshi et al. [31] and Mannan et al. [32] across focus settings. Both baseline methods employ constrained least squares optimization, penalizing deviations between patches in predicted and captured data. Mannan et al. include an additional penalty over filtered versions of the predicted and captured data (e.g., to facilitate reconstructing edges).  Fig. 3 compares the ground truth PSFs to those recovered by our method and the two baselines.
Qualitatively, our lens blur field recovers PSFs closer to the ground truth: they have smooth variations with far fewer of the high-frequency artifacts that appear in other methods as shown in  Fig. 3. We find that the proposed method also significantly outperforms the recovered PSFs in terms of quantitative accuracy, as evaluated using root-mean-square error (RMSE), indicated for each PSF displayed in  Fig. 3.
We also evaluate how well the predicted PSFs estimate the blurry calibration patterns and report the peak signal-to-noise ratio, RMSE, and structural similarity between the predicted and ground truth blurry patterns. To assess performance at withheld focal stack positions, we query our MLP at the corresponding focus setting. For the other methods, we perform linear interpolation between the nearest optimized PSFs. We find that our method outperforms the baselines when evaluating on the validation lens positions (i.e., positions not seen during training or optimized explicitly by the baselines) as shown in Tab. I. The image reconstruction metric is less sensitive to the high-frequency artifacts in the PSF obseved in Joshi et al.’s method [31], and so their approach performs roughly on par with our method in terms of overfitting to the blurry image. Supplement Sec. S-IV contains additional comparisons with neural deconvolution baselines. Results in Tab. I also show that directly obtaining PSF predictions at validation lens positions from the MLP (using the MLP’s interpolation) outperforms a baseline that obtains PSFs predictions only at training lens positions and then uses linear interpolation to estimate the validation lens positions.
Training Lens Positions Validation Lens Positions
Proposed Joshi et al. [31] Mannan & Langer [32] Proposed Joshi et al. [31] Mannan & Langer [32] Lin. Interpolation + Proposed
Training Patterns PSNR $\uparrow$ $32.109\pm 1.201$ $31.829\pm 1.016$ $15.420\pm 0.381$ $30.462\pm 3.850$ $30.797\pm 3.172$ $15.291\pm 0.513$ $29.394\pm 3.607$
SSIM $\uparrow$ $0.929\pm 0.072$ $0.950\pm 0.071$ $0.601\pm 0.021$ $0.924\pm 0.066$ $0.944\pm 0.066$ $0.600\pm 0.024$ $0.898\pm 0.077$
RMSE $\downarrow$ $0.022\pm 0.004$ $0.025\pm 0.003$ $0.166\pm 0.010$ $0.031\pm 0.023$ $0.031\pm 0.013$ $0.169\pm 0.013$ $0.034\pm 0.022$
Validation Patterns PSNR $\uparrow$ $32.051\pm 1.179$ $31.981\pm 1.277$ $16.753\pm 3.189$ $30.396\pm 3.840$ $28.964\pm 4.493$ $16.400\pm 3.086$ $29.351\pm 3.651$
SSIM $\uparrow$ $0.924\pm 0.078$ $0.916\pm 0.103$ $0.461\pm 0.290$ $0.919\pm 0.072$ $0.874\pm 0.100$ $0.452\pm 0.278$ $0.887\pm 0.091$
RMSE $\downarrow$ $0.022\pm 0.004$ $0.022\pm 0.003$ $0.139\pm 0.079$ $0.031\pm 0.023$ $0.038\pm 0.029$ $0.145\pm 0.077$ $0.034\pm 0.023$
TABLE I: Evaluation of reconstruction accuracy on synthetic noise patterns. The proposed approach quantitatively outperforms non-parametric, optimization-based methods in terms of peak signal-to-noise ratio (PSNR) and root-mean-square error (RMSE). Validation patterns have the same frequencies as the training set. MLP interpolation outperforms linear interpolation (denoted as “Lin. Interpolation + Proposed”), where PSFs at validation lens positions are obtained by linearly interpolating the MLP predictions at the training lens positions. The difference is most pronounced near the in-focus position because uniform diopter sampling yields non-uniform metric distance sampling, and PSF size changes become more significant due to a derivative sign change near focus. MLP interpolation effectively handles the resulting complex, nonlinear defocus blur radius model derived from the thin lens equation. Additional evaluations are included in Supplement Sec. S-IV.
Finally, we also validate our approach using experimental data, as presented in Tab. II. We blur a ground-truth calibration image with our learned blur fields and compare the output to real-world blurry captures. Metrics are computed on both training and a held-out validation dataset, the latter containing calibration patterns not seen during training. Our method also generalizes well to the validation dataset.
PSNR $\uparrow$ SSIM $\uparrow$
Training Dataset iPhone 14 Pro Device 0 $36.969\pm 1.926$ $0.976\pm 0.0158$
iPhone 14 Pro Device 1 $38.715\pm 1.806$ $0.977\pm 0.0157$
iPhone 12 Pro Device 0 $36.668\pm 1.587$ $0.972\pm 0.0192$
iPhone 12 Pro Device 1 $36.789\pm 1.593$ $0.959\pm 0.0369$
Pixel 4 $29.940\pm 1.635$ $0.964\pm 0.0201$
Canon EF 14mm f/2.8L $37.276\pm 1.937$ $0.996\pm 0.0028$
Canon EF 24-70mm f/2.8L $32.293\pm 0.725$ $0.985\pm 0.0476$
Canon EF 50mm f/1.2L $31.068\pm 0.411$ $0.987\pm 0.0225$
Canon EF 50mm f/1.4 $33.286\pm 0.301$ $0.991\pm 0.0231$
Validation Dataset iPhone 14 Pro Device 0 $36.736\pm 1.949$ $0.973\pm 0.0196$
iPhone 14 Pro Device 1 $38.232\pm 1.814$ $0.976\pm 0.0175$
iPhone 12 Pro Device 0 $32.605\pm 0.650$ $0.930\pm 0.0303$
iPhone 12 Pro Device 1 $31.046\pm 0.804$ $0.904\pm 0.0370$
Pixel 4 $24.559\pm 1.974$ $0.916\pm 0.0260$
Canon EF 14mm f/2.8L $37.273\pm 1.919$ $0.996\pm 0.0028$
Canon EF 24-70mm f/2.8L $31.531\pm 0.633$ $0.982\pm 0.0484$
Canon EF 50mm f/1.2L $31.134\pm 0.399$ $0.987\pm 0.0216$
Canon EF 50mm f/1.4 $33.160\pm 0.302$ $0.991\pm 0.0225$
TABLE II: Models for each device are validated by comparing artificially blurred images to blurry captured images, for both training and unseen validation datasets. In general, lenses across both the training and validation sets collectively exhibit comparable performance. Our method also produces consistent performance across training and validation sets for most devices.
IV-B Comparison to Prior Methods
The proposed method outperforms prior methods for dense stacks as shown in Tab. III. While a direct comparison of runtime and memory for a single PSF favors prior work, our method scales better with the number of images—a requirement for accurate calibration. For the example shown in Tab. III, consider optimizing a 5D PSF for a 12 MP (4032x3024) iPhone 12 Pro; if we optimize blur kernels at 75x100 sensor positions (i.e., assume each blur kernel is constant within a 40 pixel area), 15 focus positions, 73x73 blur kernel samples, and 4 RGGB color channels, then Joshi et al. (2008) would require 34 days for optimization, Mannan and Langer (2016) would require 1945 days, and our approach requires 14 hours. Moreover storing this discretely sampled 5D PSF requires 9.3 GB whereas our MLP requires 19.1 MB. Prior methods were also observed to be more noise-sensitive, often produce estimations with high-frequency artifacts, and deteriorates for large PSFs (see Fig. 3 and Supplement Sec. S-IV).
Optimization Time Memory Usage
Proposed Single PSF N/A N/A
5D PSF 14 hours 19.1MB
Joshi et al. [31] Single PSF 6.5 sec 39KB
5D PSF 34 days 9.3GB
Mannan & Langer [32] Single PSF 373.4 sec 39KB
5D PSF 1945 days 9.3GB
TABLE III: Optimization time and memory requirements for a single (73x73) blur kernel and (75x100x15x73x73) 5D PSF, where we consider optimizing a 5D PSF for a 12 MP (4032x3024) iPhone 12 Pro. If we optimize blur kernels at 75x100 sensor positions (i.e., assume each blur kernel is constant within a 40 pixel area), 15 focus positions, 73x73 blur kernel samples, and 4 RGGB color channels.
IV-C Comparing iPhone 12 Pro and 14 Pro Lens Blur Fields
Our method can resolve the differences in optical behavior between the wide-angle cameras on two separate iPhone 12 Pro devices (Supplement Sec. S-II) and two separate iPhone 14 Pro devices (Fig. 6). To demonstrate this, we conduct two types of repeatability experiments.
First, we assesses the variation in estimated blur fields for a single camera with successive captures, where the only difference is the closing and re-launching the acquisition app, while camera placement remains constant. This first type of evaluation measures the sensitivity of our method to device-centric parameters, such as focus-position repeatability or repeatability of camera software settings. Second, we evaluate the variation in estimated blur fields after closing the acquisition app, unmounting, and remounting the phone, which assesses repeatability of the entire acquisition procedure.
For both experiments, we use each device to capture 4 separate focal stack datasets and train a separate model for each repeatablity investigation (i.e., a total of 8 datasets per iPhone). Fig. 6 shows recovered 5D lens blur fields and 2D PSF slices recovered for each device, averaged across the four trials. We find that the standard deviation of estimated PSF values across trials is significantly smaller than the RMS difference between the mean PSFs of each device. Moreover, we see clear differences between the PSFs in the Supplement, especially near the PSF boundaries. Our method is sufficiently sensitive to uncover such cross-device differences without specialized hardware. Supplement Sec. S-II contains additional details on these repeatability experiments.
IV-D Visualizing xy-slices of the iPhone 12 Pro Blur Field
In the supplementary video, we visualize xy-slices of the blur fields from two different iPhone 12 Pro devices of the same make and model, showcasing the high spatial frequency content and variations across the sensor plane. This analysis highlights the challenges of capturing and analyzing such complexity using local deconvolution on a sparse patch grid.
IV-E Comparing SLR Lens Blur Fields
Our methodology demonstrates robustness against dismounting and remounting of SLR lenses, indicating negligible optical axis misalignment between lens removals. Repeatability experiments were conducted using a Canon EF 50mm f/1.4 lens for separate trials with and without removing and re-attaching the lenses (see Fig. 6). In each experiment, we captured five focal stack datasets and trained individual models for each, resulting in a total of ten datasets. Comparing the blur fields between remount and without remount experiments showed insignificant differences.
IV-F Reconstructing Dual-Pixel Lens Blur Fields
Our proposed method enables reconstructing lens blur fields for left- and right-looking pixels on dual-pixel cameras, such as the Google Pixel 4 (shown in  Fig. 1). We observe a complex dependency of the PSF on image plane location and focus setting, with the observed blur patterns departing radically from simpler parametric models for dual-pixel PSFs [10]. Notably, we find that PSF centers of mass do not always follow the expected left-right orientations, and PSFs are significantly skewed at locations near the periphery of the image sensor.
IV-G Additional SLR Lens Blur Fields
We recover lens blur fields for a variety of Canon SLR lenses of quality and type, shown in Fig. 7. Our method is sensitive enough to recover various properties of the PSFs, including the octagonal shape of the 50 mm Canon f/1.4 lens (due to the aperture blades), and the irregular shape of the Canon 24–70 mm f/2.8L zoom lens (evaluated at 50 mm focal setting), which is most likely due to vignetting. The Canon EF 14 mm f/2.8L and Canon EF 50 mm f/1.2L prime lenses show smooth, round defocus blur with minimal spatial variation.
IV-H 6D Lens Blur Fields
A 6D blur field extends a 5D blur field by including an additional dimension, $d$, for target distance. We demonstrate recovery of the 6D lens blur field for an iPhone 12 Pro’s wide lens in Fig. 4, illustrating our method’s adaptability to increased spatial dimensions. We see less variation in the PSF over target distance dimension than in other dimensions, such as focus position. This behavior is consistent with that of a thin lens, where variations in target distance can be largely predicted from a 5D blur field.
Figure 4: 6D lens blur field calibrated for an iPhone 12 Pro wide lens and the capture setup for 6D blur fields. We show 5D slices of the 6D function, which vary over focus distance (top left) and target distance (top right), ranging from 20.5 cm to 41.5 cm. While we observe small changes in the PSF over target distance, the most dramatic changes occur over the focus dimension. This is expected as smartphone cameras have a small apertures, demonstrating that a 5D blur field may be a sufficient blur characterization for certain applications. The 6D capture setup comprises a monitor, a phone stand, and a translating stage that controls positioning along the target distance dimension.
IV-I Partial-Field-of-View Blur Field Acquisition
Our method successfully acquires a full-sensor blur field, even in scenarios where the display spans a partial field of view of the sensor.  Fig. 5 shows an experiment where one blur field was trained on two sets of captures, with the patterns partially observed in each set.
Figure 5: Demonstration of partial-field-of-view capture using an iPhone 12 Pro wide lens. Two sets of focal stacks are captured, spanning roughly 55% of the image plane’s width. We follow the same post-capture processing scheme for each set separately and train the network on both datasets. The reconstructed blur field yields similar results to the full-field-of-view (FOV) reconstruction (bottom right).
V Rendering with Blur Fields
Lens blur fields can be used as lightweight computational proxies for underlying device optics. We demonstrate this principle with a few examples where we use it to render synthetic scenes with device-specific blur.
V-A Lens Blur Field Dataset
To facilitate applications of lens blur fields in computational photography and rendering, we make available a dataset of lens blur fields calibrated across a range of smartphone cameras and SLR lenses. The dataset includes all lens blur fields described thus far (a complete listing of cameras and lenses included in the dataset is provided in Supplement Tab. S-III).
V-B Resolution Chart Experiment
A comparison of captures and estimates rendered with the calibrated lens blur field of the iPhone 12 Pro visualized in  Fig. 5 is presented in the Supplement. We capture a focal stack of a resolution chart and the captured images are compared with image restorations, showing overall agreement across the focus range (see Fig. 8).
V-C Improved Image Restoration with Blur Fields Compared to Seidel Model
We compare the proposed lens blur fields with the parametric Seidel model by evaluating their performance for image restoration on a real-world capture of a Siemens Star test pattern using an iPhone 12 Pro wide camera (Tab. II). After constructing the lens blur field, we follow Kohli et al. [80] to fit a Seidel model—with spherical, coma, astigmatism, field curvature, and defocus coefficients—to the PSFs recovered by the blur field by randomly sampling 32 image plane locations per lens position. The Seidel coefficients are optimized via gradient descent to match the lens blur field PSFs at the sampled locations using 300 iterations with a learning rate of $0.01$, the default parameters in Kohli et al. [80]. For image restoration, we use the same optimization approach as for PSF estimation (Sec. II-B). We solve for a restored image that, when convolved with the respective blurs, produces a synthetic blurry image. We then compute a loss between this synthetic image and the real-life blurry image.
We compare the resulting restorations and find that blur field PSFs yield superior results. Qualitative results in Fig. 8 demonstrate that the lens blur field better preserves high-frequency details in the Siemens Star pattern, reconstructing longer portions of the star’s arms than do restorations using PSFs from the Seidel model. A direct comparison with the ground truth Siemens Star in Tab. IV shows that restorations using blur field PSFs are superior. Furthermore, the synthetic blurry image generated by convolving the restored Siemens Star with blur field PSFs more closely matches the real-life image than the synthetic image generated using Seidel PSFs.
Restoration Results Siemens Star Restoration Synthetic Blurry Image
Red Green 1 Green 2 Blue
Proposed PSNR $\uparrow$ 8.900 20.335 20.392 21.479 17.700
SSIM $\uparrow$ 0.293 0.747 0.780 0.786 0.600
RMSE $\downarrow$ 0.353 0.050 0.054 0.047 0.059
Seidel Coefficients PSNR $\uparrow$ 7.415 20.184 19.660 20.528 17.754
SSIM $\uparrow$ 0.127 0.534 0.588 0.595 0.515
RMSE $\downarrow$ 0.421 0.051 0.059 0.053 0.058
TABLE IV: Restoration results for both the estimated Siemens Star and the synthetic blurry image obtained by convolving the estimated Star with either the blur fields PSF or the Seidel PSF. Both the restored Siemens Star and the synthetic blurry image show that the blur field PSFs recover real-life effects more effectively than the Seidel model.
V-D Rendering Scenes with Blur
To facilitate synthetic rendering, we create a Blender [81] postprocessing add-on that allows one to take advantage of our pre-trained lens blur field models as part of their rendering pipeline. We demonstrate this technique on synthetic scenes shown in Supplement Sec. S-V, rendered with the occlusion-aware model of Ikoma et al. [82] (which itself is agnostic to whether Blender or another engine is used). Different lens blur fields result in a diverse set of bokeh patterns and we show an additional result on captured data in Supplement Sec. S-V.
VI Concluding Remarks
Although billions of people carry around smartphone cameras in their pockets, the precise optical behavior of individual devices is inaccessible, or entirely unknown, to the owner. Proprietary databases exist which characterize optical behavior generally for a make and model of a lens, [^1] For example, DxO has one such database described at https://www.dxo.com/technology/how-dxo-corrects-lens-flaws/. but calibrating device-specific high-dimensional models of optical blur with commodity hardware has largely been an intractable problem. We take steps towards democratizing accurate optical characterization via the lens blur field, and provide a robust, lightweight method by which it can be acquired.
We are excited about several avenues of future work. For example, given that we find significant differences between lens blur fields from devices of the same make and model, how distinct are these differences across a wider cross-section of device types and instances? What can we learn about the devices from these aberrations? Additionally, we wish to explore how the lens blur field can be extended to even higher dimensions (e.g., light fields, hyperspectral data, polarization imaging, etc.).
It may also be possible to improve the optimization procedure by incorporating pre-processing steps (such as geometric alignment and auto-exposure) into the main optimization loop. Tackling the blind deconvolution problem would also enable optical characterization in entirely unstructured environments.
Finally, neural blur fields may open up new applications as proxy functions for emulating optical systems. Here, we envision uses for photorealistic rendering and synthetic depth of field for mobile photography.
Figure 6: (A) Comparison between calibrated PSFs from the wide cameras of two iPhone 14 Pro devices. We estimate PSFs using four separate sets of calibration measurements captured independently for each device, for each repeatability investigation. Visualizing slices of the mean estimated 5D PSF function for each device (top) reveals broad agreement in the PSF structure; however, significant variations can be seen upon close inspection (insets, bottom). While imperfections in the calibration and capture process result in differences between PSF estimates for the same device (see standard deviation of PSFs, bottom left and right), this effect is generally weaker than PSF differences between devices (bottom center, shown as root-mean-square deviation), suggesting that each device has a specific blur signature. (B) Comparison between static and dynamic experiments for a Canon EF 50mm f/1.4 lens. We capture five separate sets of measurements for each repeatbility experiment (with and without remount) and visualize the mean and standard deviations for each experiment with the same color scale. We observe little differences between experiment, indicating that remounting the lens does not alter the blur field of the lens.
Figure 7: Reconstructed lens blur fields for a collection of Canon lenses. Our method recovers interesting features in the PSFs of the Canon EF 50 mm f/1.4 lens and the Canon EF 24–70 mm zoom lens (evaluated at the 50 mm focal position). For example, the f/1.4 lens has octagonal PSFs from the aperture blades, and the PSF of the zoom lens is distorted, perhaps due vignetting effects. The prime 14 mm f/2.8 and 50 mm f/1.2 lenses have round, smooth PSFs.
Figure 8: (A) Comparison of real captures and synthetically generated renders using a learned blur field from the iPhone 12 Pro wide lens, and (B) restoration results from these captures via non-blind deconvolution using our blur field PSFs versus a Seidel model fitted to our PSFs. A resolution chart (Edmund Optics I3A/ISO 12233 Digital Camera Resolution Test Chart, #58-940) is positioned at a distance of 36 cm from an iPhone 12 Pro. The chart is scanned at a resolution of 600 DPI and subsequently blurred at the different focus settings indicated, using the previously-captured blur field (visualized in  Fig. 4). The 2D PSF at the center of each region is used for blurring and shown as insets. There is overall agreement between the captured and estimated images across the entire focus range. In the restoration results, we perform deconvolution using the proposed method. Our forward model convolved the ground truth estimate with PSFs to create a synthetic image, which is then compared to the actual capture to compute the loss. The blurry capture is boxed in orange, and the PSFs from blur fields (boxed in blue) and Seidel are shown for each color channel. While the Seidel PSFs are comparable in size to the blur field PSFs, they do not capture real-life effects as effectively. The lens blur field preserves high-frequency details in the Siemens Star arms better than the Seidel PSF, as seen in the yellow-boxed region. The synthetic blurry image generated with blur field PSFs preserves more detail and more closely resembles the captured blurry image than that generated with the Seidel PSFs.
Acknowledgment
DBL and KNK acknowledge support of NSERC under the RGPIN, RTI, and USRA programs. DBL also acknowledges support from the Canada Foundation for Innovation and the Ontario Research Fund.